# ADHint: Adaptive Hints with Difficulty Priors for Reinforcement Learning

### v1

@@ 4.3 Advantage Estimation with Rollout Difficulty Posterior
Specifically, for each query $q$, we define the difficulty score of naive-rollouts in Eq. 1. And similarly define the difficulty score of hint-rollouts as
$\text{Diff}_{\text{H}}\;=\;1-\mathcal{M}(n+1:)\;=\;1-\operatorname{mean}(r_{n+1},\dots,r_{n+m}).$ (6)
@@
$\text{Diff}_{i}\;=\;\begin{cases}\text{Diff}_{\text{N}},&1\leq i\leq n,\\ \text{Diff}_{\text{H}},&n{+}1\leq i\leq n{+}m.\end{cases}$
Let $\hat{A}_{i}$ denote the relative advantage of rollout $o_{i}$ estimated by pooling all n+mn{+}m trajectories. We define $s_{i}$ as a sign indicator:
$s_{i}\;=\;\begin{cases}1,&\hat{A}_{i}>0,\\ -1,&\hat{A}_{i}\leq 0.\end{cases}$
Finally, the relative advantages in ADHint are estimated as
$\tilde{A}_{i}\;=\;\left(\frac{\mathcal{M}(1:n+m)}{1-\text{Diff}_{i}}\right)^{s_{i}}\hat{A}_{i}.$ (7)
@@ 4.4 Overall Gradient Computation
$\nabla_{\theta}\mathcal{J}^{*}_{\text{ADHint}}(\pi_{\theta})\Big|_{q,o_{i}}=\frac{1}{G\cdot|o_{i}|}$ (8)
$\sum_{t=1}^{|o_{i}|}k_{i,t}\cdot\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\cdot\tilde{A}_{i}\cdot\nabla_{\theta}\log(\pi_{\theta}(o_{i,t}|q,o_{i,<t})).$
Notably, to avoid inaccurate difficulty prior evaluation caused by formatting errors, we warm up the policy with GRPO for the first five steps so that the model follows the required response format (Figure 5(f)).
(a) Training reward
(b) Policy entropy
(c) Gradient norm
(d) Response length
(e) Clipping ratio of long responses
(f) Format reward
@@ 5 Experiments
Table 1: Performance comparison of ADHint and baselines on Qwen2.5-VL across six benchmarks spanning three domains. For each benchmark, we report pass@1 (the first value) and avg@8 (the second value). The upper block uses Qwen2.5-VL-3B as the base MLLM, while the lower block uses Qwen2.5-VL-7B. The best and runner-up results are highlighted in bold and underlined, respectively.
5.1 Settings
Training data.
We conduct extensive experiments across both MLLMs and LLMs to validate ADHint’s effectiveness. For MLLMs, we construct our training data from the ViRL39K dataset [wang2025vl], which spans multiple domains, including STEM, chart reasoning, document reasoning, and spatial reasoning. We use Qwen3-VL-30B-A3B-Thinking and Qwen3-VL-235B-A22B-Thinking [bai2025qwen3vltechnicalreport] to sample hints, and remove examples that are considered excessively simple for Qwen2.5-VL-7B-Instruct [bai2025qwen2], resulting in a final set of 17K training queries, termed ViRL-Hint17k. For LLMs, we use 21K examples sampled from public training data [yan2025learning], with hints generated by DeepSeek-R1 [guo2025deepseek]. More details are provided in Appendix C.1.
Benchmarks.
For MLLM evaluation, we use six benchmarks spanning diverse domains, including MathVista [lu2023mathvista], MathVerse [zhang2024mathverse], and We-Math [qiao2024we] for math reasoning; MMMU [yue2024mmmu] and MMMU-Pro [yue2024mmmupro] for multi-discipline reasoning; and LogicVista [xiao2024logicvista] for logical reasoning. For LLM evaluation, we use AIME24, AIME25, AMC [li2024numinamath], Minerva [lewkowycz2022solving], OlympiadBench [he2024olympiadbench], and MATH500 [hendrycks2measuring]. Further details are provided in Appendix C.2.
Evaluation Metrics.
For MLLM evaluation, we report pass@1 and avg@8 for each benchmark. Pass@1 measures the accuracy of a single rollout per sample under greedy decoding, reflecting the model’s out-of-distribution reasoning generalization. Avg@8 computes the average accuracy over eight rollouts per sample at temperature 1.0, reflecting the model’s robustness and mastery of the underlying knowledge. For LLM evaluation, we report either pass@1 or avg@32 depending on the benchmark to mitigate potential variance introduced by the small sizes of some datasets.
Implementation Details.
For MLLM training, we use Qwen2.5-VL-7B-Instruct, Qwen2.5-VL-3B-Instruct, and Qwen3-VL-8B-Instruct, and train them for 250 steps (about 2 epochs). We set the learning rate to $1\times 10^{-6}$, the rollout batch size to 128, and the maximum generation length to 8K. For LLM training, we use Qwen2.5-Math-7B [yang2024qwen2]. Across all experiments, we set $w_{\text{max}}=0.2$ and $w_{\text{min}}=0.0$, the sampling radius of the noise term $\sigma$ to $0.01$, and the threshold in CGM to $\alpha=0.5$. We perform all training on 4 nodes, each equipped with 8 H100 GPUs. Further details are provided in Appendix C.3.
Baselines.
We compare ADHint with several baselines: (1) GRPO; (2) SFT, trained for two epochs on the full training set; (3) SFT $\rightarrow$ RL, which applies two epochs of SFT followed by GRPO; (4) LUFFY [yan2025learning], which performs on-policy and off-policy updates jointly at each step; (5) StepHint [zhang2025stephint], which mixes hint-rollouts with multiple hint ratios, naive-rollouts, and ground-truth trajectories into a group for update; (6) HintGRPO [Huang_2025_ICCV], which progressively increases the hint ratio for the hardest samples to obtain positive trajectories; and (7) GHPO [liu2025ghpo], which treats hints as part of the query and uses multiple hint ratios for the hardest samples. We note that several baselines crash during training, and for these methods, we report results from the last checkpoint available before the crash. More detailed information about the baselines and their reimplementation is provided in Appendix C.4 and Appendix C.5, respectively.
Bench. Metric GRPO HintGRPO ADHint $\Delta$
MathVista p@1 74.6 64.9 74.0 0.6
a@8 78.8 64.4 79.6 0.8
MathVerse p@1 63.1 64.2 75.2 12.1
a@8 79.0 68.7 80.1 1.1
We-Math p@1 73.1 68.6 76.4 3.3
a@8 82.6 71.7 85.6 3.0
MMMU p@1 61.0 47.2 64.4 3.4
a@8 71.7 55.5 70.1 1.6
MMMU-Pro p@1 46.5 38.4 51.8 5.3
a@8 57.9 46.0 57.9 0.0
LogicVista p@1 49.4 43.1 56.3 6.9
a@8 63.0 51.9 64.9 1.9
Avg. p@1 61.3 54.4 66.4 5.1
a@8 72.2 59.7 73.0 0.8
Table 2: Performance comparison of ADHint and baselines on Qwen3-VL-8B-Instruct. We report p@1 (pass@1) and a@8 (avg@8) for each method. The last column shows the absolute performance differences of ADHint relative to GRPO, where green indicates improvement and red indicates degradation.
Figure 6: Performance comparison of ADHint and GRPO on Qwen2.5-Math-7B. For AIME and AMC, we report avg@32 due to their small sizes, and we report pass@1 for the remaining benchmarks.
@@ 5.1 Settings
Table 4: Effect of the maximum hint ratio $w_{\text{max}}$. For each benchmark, we report pass@1 (the first value) and avg@8 (the second value).
@@
Table 5: Effect of the threshold $\alpha$ in CGM. For each benchmark, we report pass@1 (the first value) and avg@8 (the second value).
5.2 Main Results
SOTA performance with Qwen2.5-VL.
In Table 5, we compare the proposed ADHint against several baselines on Qwen2.5-VL-7B and Qwen2.5-VL-3B. ADHint delivers consistent performance gains at both model scales. At the 7B scale, ADHint improves pass@1 by 2.1% and avg@8 by 2.0% over GRPO. Relative to recent methods, it delivers larger gains of 6.8% in pass@1 and 11.1% in avg@8. ADHint also performs well at the 3B scale, improving pass@1 by 2.3% and avg@8 by 2.1% over GRPO, and by 9.7% and 10.5% over recent methods, respectively. These results confirm ADHint’s advanced multimodal reasoning ability and suggest that integrating difficulty into both the hint-ratio schedule and relative-advantage estimation provides a feasible solution for hint-based on-policy learning on challenging multimodal tasks. In Appendix D, we provide a more in-depth comparative analysis of the performance differences between ADHint and the baselines.
Performance with Stronger Backbone.
Table 2 presents a comparison between ADHint and several baselines on Qwen3-VL-8B-Instruct, a recent strong backbone that achieves state-of-the-art performance among models of similar scale across multiple multimodal reasoning benchmarks. Compared with GRPO, ADHint achieves competitive performance on most metrics, with absolute gains of 5.1% in pass@1 and 0.8% in avg@8. Notably, ADHint achieves gains of over 5.0% in pass@1 on MathVerse, MMMU-Pro, and LogicVista, which cover challenging problems from three different domains. These results highlight the considerable promise of ADHint for improving general multimodal capabilities. As future work, we will collect more challenging datasets tailored to the Qwen3-VL series and leverage hints from stronger models to further unlock ADHint’s potential.
Performance with LLM.
In Figure 6, we further evaluate the effectiveness of ADHint on text-only reasoning. ADHint improves the average accuracy by 2.3% over GRPO across the six math reasoning benchmarks. On OlympiadBench, which covers Olympiad-level questions, and AIME25, which consists of challenging problems from the 2025 AIME, ADHint surpasses GRPO by 5.7% and 2.4%, respectively. These results demonstrate the strong generalization of ADHint, which broadens the capability boundaries of models across different modalities.
5.3 Training Dynamics
Figure 5 presents the training dynamics of several key metrics. Figure 5(a) shows that ADHint consistently maintains rewards within a moderate regime, yielding low-variance update signals. Figures 5(b) and 5(c) indicate that ADHint maintains exploration within a healthy entropy range with more stable gradient updates. Figures 5(d) and 5(e) show that ADHint steadily increases response length, while the proportion of overlength responses does not increase, which usually indicates repetition or garbled outputs. These trends suggest that the policy model continuously acquires effective long-term thinking capabilities. Taken together, these observations indicate that ADHint stably and adaptively learns reasoning ability from off-policy hints.
5.4 Ablation Study
Table 3 reports the ablation results of four critical components in ADHint. Specifically, “w/o AH-SDP” replaces our adaptive hint-ratio schedule with an annealing hint-ratio strategy, while “w/o AE-RDP” pools hint-rollouts and naive-rollouts into a single group for advantage estimation. Removing any component results in performance degradation on both pass@1 and avg@8, underscoring the importance of our design. The difficulty-aware design in hint-ratio schedule and relative-advantage estimation ensures efficient learning that balances exploration and imitation. The gradient modulation mechanisms prevent biased and destructive updates, ensuring healthy policy learning.

### v2

@@ 4.3 Advantage Estimation with Rollout Difficulty Posterior
Specifically, for each query $q$, we define the difficulty score of naive-rollouts in Eq. 1, and similarly define the difficulty score of hint-rollouts as
$\mathrm{Diff}_{\mathrm{H}}=1-\mathcal{M}(n+1:)=1-\operatorname{mean}(r_{n+1},\dots,r_{n+m}).$ (6)
@@
$\mathrm{Diff}_{i}=\mathrm{Diff}_{\mathrm{N}}\text{ if }1\leq i\leq n,\text{ and }\mathrm{Diff}_{\mathrm{H}}\text{ if }n+1\leq i\leq n+m.$
Let $\hat{A}_{i}$ denote the relative advantage of rollout $o_{i}$, estimated by pooling all $n+m$ trajectories. We define a sign indicator $s_{i}=\operatorname{sgn}(\hat{A}_{i})$, where $s_{i}=-1$ when $\hat{A}_{i}\leq 0$. Finally, the relative advantages in ADHint are estimated as
$\tilde{A}_{i}=\left(\frac{\mathcal{M}(1:n+m)}{1-\mathrm{Diff}_{i}}\right)^{s_{i}}\hat{A}_{i}.$ (7)
@@ 4.4 Overall Gradient Computation
$\nabla_{\theta}\mathcal{J}^{*}_{\text{ADHint}}(\pi_{\theta})\Big|_{q,o_{i}}=\frac{1}{G\cdot|o_{i}|}\sum_{t=1}^{|o_{i}|}k_{i,t}\cdot\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\cdot\tilde{A}_{i}\cdot\nabla_{\theta}\log(\pi_{\theta}(o_{i,t}|q,o_{i,<t})).$
Notably, to avoid inaccurate difficulty prior evaluations caused by formatting errors, we warm up the policy model with GRPO for the first five steps to ensure the model follows the required response format (Figure 4(h)).
@@ 5 Experiments
5.1 Settings
We conduct experiments across multiple modalities, model scales, model families, and domains to validate ADHint’s effectiveness. For common MLLM reasoning tasks (e.g., math and chart reasoning), we evaluate ADHint on Qwen2.5-VL-7B-Instruct [bai2025qwen2], Qwen3-VL-8B-Instruct [bai2025qwen3vltechnicalreport], Qwen2.5-VL-3B-Instruct, Qwen3-VL-4B-Instruct, and MiMo-VL-7B-SFT-2508 [coreteam2025mimovltechnicalreport]. We additionally evaluate ADHint on Medical VQA to assess cross-domain generalization, using the Qwen2.5-VL-7B-Instruct backbone. For LLM math reasoning, we use Qwen2.5-Math-7B [yang2024qwen2], following the data and experimental settings of GHPO [liu2025ghpo].
Training data.
For MLLMs, we construct our training data from the ViRL39K dataset [wang2025vl], which spans multiple domains, including STEM, chart reasoning, document reasoning, and spatial reasoning. We use Qwen3-VL-30B-A3B-Thinking and Qwen3-VL-235B-A22B-Thinking to sample hints, and remove examples that are considered excessively simple for Qwen2.5-VL-7B-Instruct, resulting in a final set of 17K training queries, termed ViRL-Hint17k. For Medical VQA, we use a 15K subset of the PMC-VQA training set [zhang2023pmc]. For LLMs, we use NuminaMath-S, a mixture of MATH [hendrycks2measuring] and NuminaMath-1.5 [hendrycksmath2021] curated by GHPO. More details are provided in Appendix D.1.
Benchmarks.
For MLLMs, we use benchmarks spanning diverse domains, including MathVista [lu2023mathvista], MathVerse [zhang2024mathverse], and We-Math [qiao2024we] for math reasoning; MMMU [yue2024mmmu] and MMMU-Pro [yue2024mmmupro] for multi-discipline reasoning; and LogicVista [xiao2024logicvista] for logical reasoning. For Medical VQA, we use the PMC-VQA test set (PMC-Test) [zhang2023pmc]. For LLMs, we evaluate on math reasoning benchmarks, including AIME [patel2024aime], AMC [li2024numinamath], Minerva [lewkowycz2022solving], OlympiadBench [he2024olympiadbench], and MATH500 [hendrycks2measuring]. More details are provided in Appendix D.2.
Evaluation Metrics.
For MLLMs, we report pass@1 and avg@8 for each benchmark [wang2025vl, wang2025perception]. Pass@1 measures the accuracy of a single rollout per sample under greedy decoding, reflecting the model’s out-of-distribution reasoning generalization. Avg@8 computes the average accuracy over eight rollouts per sample at temperature 1.0, reflecting the model’s robustness and mastery of the underlying knowledge. For Medical VQA, we report pass@1 [zhu2025toward]. For LLMs, we report either pass@1 or avg@32 depending on the benchmark to mitigate potential variance introduced by the small sizes of some datasets [liu2025ghpo].
Implementation Details.
For MLLMs, we train for 250 steps (about 2 epochs) with a learning rate of $1\times 10^{-6}$. We set the rollout batch size to 128 and the maximum generation length to 8K. For LLMs, we train for 5 epochs with an initial learning rate of $1\times 10^{-6}$, decayed to zero using a cosine schedule with a warm-up phase covering 10% of the total global steps. We set the maximum generation length to 2K. We set $w_{\text{max}}$ to 0.2 for MLLMs and 0.4 for LLMs, and use $\alpha=0.5$ as the CGM threshold for both. All training is performed on 4 nodes, each equipped with 8 H100 GPUs. Further details are provided in Appendix D.3.
Baselines.
We compare ADHint with several baselines: (1) GRPO; (2) SFT, trained for two epochs on the full training set; (3) SFT $\rightarrow$ RL, which applies two epochs of SFT followed by GRPO; (4) LUFFY [yan2025learning], which performs on-policy and off-policy updates jointly at each step; (5) StepHint [zhang2025stephint], which mixes hint-rollouts with multiple hint ratios, naive-rollouts, and ground-truth trajectories into a group for update; (6) HintGRPO [Huang_2025_ICCV], which progressively increases the hint ratio for the hardest samples to obtain positive trajectories; and (7) GHPO [liu2025ghpo], which treats hints as part of the query and uses multiple hint ratios for the hardest samples. We note that several baselines crash during training, and for these methods, we report results from the last checkpoint available before the crash. More detailed information about the baselines and their reimplementation is provided in Appendix D.4 and Appendix D.5, respectively.
5.2 Main Results
Table 1: Comparative performance on Qwen2.5-VL models. For each benchmark, we report pass@1 (the first value) and avg@8 (the second value). The upper block uses Qwen2.5-VL-3B as the base MLLM, while the lower block uses Qwen2.5-VL-7B. The best and runner-up results are highlighted in bold and underlined, respectively.
@@ 5.2 Main Results
Table 2: Math reasoning performance on Qwen2.5-Math-7B. We report avg@32 on AIME and pass@1 on all other benchmarks. Results are taken from GHPO [liu2025ghpo].
Method AIME AMC Minerva Math Olympiad Avg.
GRPO [shao2024deepseekmath] 27.0 62.5 34.6 81.0 44.8 50.0
LUFFY [yan2025learning] 31.1 67.5 37.9 81.8 44.0 52.5
StepHint [zhang2025stephint] 29.6 61.4 36.4 81.0 40.4 49.8
HintGRPO [Huang_2025_ICCV] 32.8 71.1 35.3 83.0 44.7 53.4
GHPO [liu2025ghpo] 32.0 70.0 38.2 82.2 45.3 53.5
[…]
Table 3: Medical VQA performance on PMC-Test using Qwen2.5-VL-7B.
Method PMC-Test [zhang2023pmc]
GRPO [shao2024deepseekmath] 53.4
LUFFY [yan2025learning] 49.4
StepHint [zhang2025stephint] 48.3
HintGRPO [Huang_2025_ICCV] 20.9
GHPO [liu2025ghpo] 52.9
[…]
SOTA performance with MLLMs.
As detailed in Tables 5.2, 3, and 5, ADHint achieves consistent gains across multiple backbones, validating its robust generalization regarding model capabilities, scales, and families. For Qwen2.5-VL models, ADHint consistently outperforms GRPO, yielding overall pass@1/avg@8 improvements of 2.1%/2.0% at the 7B scale and 2.3%/2.1% at 3B. Against recent hint-based methods, it delivers larger gains of 6.8%/11.1% (7B) and 9.7%/10.5% (3B). Furthermore, for Qwen3-VL models, ADHint maintains substantial effectiveness, boosting pass@1/avg@8 by 5.1%/0.8% at 8B and 4.1%/3.0% at 4B over the baselines. Beyond the Qwen series, ADHint also achieves a 2.0%/2.2% improvement over the baselines on MiMo-VL-7B-SFT-2508. We observe that several baselines perform worse than vanilla GRPO, which is consistent with our motivation and the training dynamics observed. Methods that use fixed, high hint ratios (e.g., GHPO, HintGRPO) can introduce excessive guidance and yield extremely high training rewards (often >0.9>0.9), but tend to overfit to the hint distribution, resulting in limited generalization and can even degrade reasoning ability without hints. Additionally, several methods (e.g., StepHint, LUFFY) ignore the imbalance in intra-group advantage estimation and the sharp gradients induced by hint tokens, leading to unstable update signals and training collapse. Moreover, our ViRL-Hint17k is more challenging, featuring longer reasoning chains (3100 tokens on average) than the relatively short baselines’ data (typically a few hundred tokens, e.g., $\sim$230 in GHPO), which further exposes these shortcomings. To address these issues, ADHint introduces novel mechanisms for hint-ratio schedule, relative-advantage estimation, and hint gradient modulation, providing a feasible solution for hint-based on-policy learning on challenging multimodal tasks. A more in-depth comparative analysis of ADHint and the baselines is provided in Appendix E.
Table 4: Comparative performance on Qwen3-VL models. For each benchmark, we report pass@1 (the first value) and avg@8 (the second value).
Method MathVista MathVerse We-Math MMMU MMMU-Pro LogicVista Avg.
[…]
GRPO [shao2024deepseekmath] 73.6/77.9 56.1/75.9 67.9/82.1 38.1/49.5 39.4/51.2 46.2/60.8 53.6/66.2
LUFFY [yan2025learning] 64.7/72.3 56.3/70.9 56.6/74.5 29.9/46.6 32.0/49.7 38.2/57.1 46.3/61.9
StepHint [zhang2025stephint] 66.2/71.2 53.7/70.2 58.2/74.0 34.7/52.5 32.8/49.4 38.0/57.0 47.3/62.4
HintGRPO [Huang_2025_ICCV] 67.1/71.9 60.3/72.8 64.6/77.6 42.0/52.2 40.1/49.4 48.0/57.2 53.7/63.5
GHPO [liu2025ghpo] 70.4/75.9 54.9/70.5 69.5/76.4 41.2/48.5 39.8/51.1 48.4/59.0 54.0/63.6
[…]
GRPO [shao2024deepseekmath] 74.6/78.8 63.1/79.0 73.1/82.6 61.0/71.7 46.5/57.9 49.4/63.0 61.3/72.2
LUFFY [yan2025learning] 54.4/53.7 28.0/18.6 26.0/18.1 21.0/26.6 21.2/27.2 14.5/21.8 27.5/27.7
StepHint [zhang2025stephint] 55.0/54.8 28.4/20.5 25.8/17.5 22.5/27.7 22.3/27.2 20.5/22.5 29.1/28.4
HintGRPO [Huang_2025_ICCV] 64.9/64.4 64.2/68.7 68.6/71.7 47.2/55.5 38.4/46.0 43.1/51.9 54.4/59.7
GHPO [liu2025ghpo] 64.9/63.2 39.7/37.7 49.9/48.8 38.8/38.5 40.4/40.0 45.1/43.6 46.5/45.3
[…]
Table 5: Comparative performance on MiMo-VL-7B-SFT-2508.
Method MathVista MathVerse We-Math MMMU MMMU-Pro LogicVista Avg.
GRPO [shao2024deepseekmath] 76.8/77.6 75.9/77.3 78.6/79.7 64.5/65.9 50.6/51.2 63.5/60.0 68.3/68.6
LUFFY [yan2025learning] 70.2/68.1 64.5/68.7 60.7/67.5 48.2/61.4 35.2/47.4 41.7/53.7 53.4/61.1
StepHint [zhang2025stephint] 72.3/67.9 64.5/67.8 63.2/66.6 36.2/43.3 37.0/47.2 43.3/54.6 52.8/57.9
HintGRPO [Huang_2025_ICCV] 80.4/79.0 79.0/77.8 81.4/79.9 57.6/56.6 53.0/52.1 59.7/60.5 68.5/67.7
GHPO [liu2025ghpo] 78.8/76.4 72.7/61.7 78.7/74.6 63.7/60.4 50.8/46.6 61.8/56.4 67.8/62.7
[…]
Performance with LLM.
Table 3 evaluates ADHint on text-only reasoning. ADHint improves average accuracy by 2.4% over baselines across math reasoning benchmarks, indicating that it can broaden model capability boundaries across different modalities. Notably, the gains are consistent across benchmarks of varying difficulty, suggesting that ADHint benefits both routine and harder reasoning problems.
Generalization to different domains.
Table 3 reports results on the Medical VQA domain, which is knowledge-intensive, more OOD for the backbone models, and substantially different from standard math reasoning. We observe that the baselines do not generalize well, whereas ADHint improves accuracy by 1.7% over GRPO. These results suggest that ADHint can be readily transferred to more challenging real-world domains.
5.3 Training Dynamics
Figure 5 presents the training dynamics of several key metrics. Figure 4(c) shows that ADHint consistently maintains rewards within a moderate regime, yielding low-variance update signals. Figures 4(d) and 4(e) indicate that ADHint maintains exploration within a healthy entropy range with more stable gradient updates. Figures 4(f) and 4(g) show that ADHint steadily increases response length, while the proportion of overlength responses does not increase, which usually indicates repetition or garbled outputs. These trends suggest that the policy model continuously acquires effective long-term thinking capabilities. Taken together, these observations indicate that ADHint stably and adaptively learns reasoning ability from off-policy hints.
(c) Training reward
(d) Policy entropy
(e) Gradient norm
(f) Response length
(g) Clipping ratio of responses
(h) Format reward
5.4 Ablation Study
Table 6 reports the ablation results of four critical components in ADHint. Specifically, “w/o AH-SDP” replaces our adaptive hint-ratio schedule with an annealing hint-ratio strategy, while “w/o AE-RDP” pools hint-rollouts and naive-rollouts into a single group for advantage estimation. Removing any component results in performance degradation on both pass@1 and avg@8, underscoring the importance of our design. The difficulty-aware design in the hint-ratio schedule and relative-advantage estimation ensures efficient learning that balances exploration and imitation. The gradient modulation mechanisms prevent biased and destructive updates, ensuring healthy policy learning.
@@ 5.4 Ablation Study
Table 7: Effect of the maximum hint ratio $w_{\text{max}}$.
@@
Table 8: Effect of the threshold $\alpha$ in CGM.